OpenStudy (anonymous):

Evaluate 1/n (summation of x where x is from 0 to n)

OpenStudy (anonymous):

\[\sum_{k=0}^n\frac{1}{k}\]?

OpenStudy (anonymous):

\[\sum_{x=0}^n\frac{1}{n}\] maybe ?

OpenStudy (anonymous):

question is not really clear if it is the second one, then you are adding up \[\frac{1}{n}+\frac{1}{n}+...+\frac{1}{n}\] with \(n+1\) terms

OpenStudy (anonymous):

No, it's just summation of x

OpenStudy (anonymous):

\[\sum_{x=0}^n\frac{1}{x}\] ??

OpenStudy (anonymous):

you can't do that